# Should We Still Pretrain Encoders with Masked Language Modeling?

### v1

@@ Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM objective or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 30 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM, achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models (from the existing LLM ecosystem), reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
@@ 1 Introduction
In this paper, we conduct a controlled study to investigate the impact of training with MLM and CLM objectives when learning textual representations. We compare models with equal architecture and size, trained on equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively on either objective, and then investigate a two-stage pretraining alternative, where training starts with the CLM objective and is followed by MLM. Additionally, given the availability of existing pretrained models, we investigate a continued pretraining scenario, where models are initialized from checkpoints trained exclusively with MLM or CLM, and then trained with MLM for a fraction of the pretraining steps. In total, we train 30 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs, amounting to 110k MI250X GPU hours. Backed by these large-scale experiments, our results show that:
@@ 2 Experimental Setup
The default training setup uses 100 billion tokens, with a per-device batch size of 12 across 192 GPUs, resulting in an effective batch size of 2,359,296 tokens. We employ a Warmup-Stable-Decay (WSD) learning rate schedule: a 2,000-step warmup phase, followed by 38,000 steps with a constant learning rate of 5e-4, [^5] The optimal learning rate was selected from the range 1e-4 to 2e-3 based on experiments using 10% of the training data. ending with a 2,000-step linear decay phase, for a total of 42,000 training steps.
@@
To ensure a fair comparison, all models are fine-tuned using a consistent protocol. Each model is trained for up to 1,000 steps or one full epoch, whichever comes first, using a batch size of 32. To account for differences in model architecture and task requirements, we perform a grid search over 6 learning rates (1e-5, 2e-5, 5e-5, 1e-4, 2e-4, and 5e-4) for each model-dataset pair, with 10% warmup followed by linear decay. The learning rate yielding the best validation performance is selected. Fine-tuning employs bidirectional attention mask with task-specific loss functions: for SC, we use cross-entropy on mean-pooled token embeddings; TC and QA are trained using token-level cross-entropy; and for IR, we rely on the InfoNCE loss (Oord et al., 2018) with in-batch negatives, using mean pooling. The entire procedure is repeated across 5 random seeds to account for training variability. Fine-tuning is conducted on the in-domain training set, except for NQ and MLDR, for which training is performed on MS MARCO.
@@
SC is assessed with accuracy, TC and QA with F1 score, and IR with NDCG@10. Results are averaged over the 5 random seeds, with 95% confidence intervals reported.
@@
Drawing reliable conclusions about model design choices typically requires scaling up experiments, as statistically sound results often require a large number of runs and sufficient training to reach minimal convergence (Hoffmann et al., 2022). We therefore carefully design our experiments to ensure robust and well-supported findings. Pretraining runs are executed across 24 nodes with 8 GPUs each (192 GPUs in total), while fine-tuning is performed on single GPUs. The total compute budget amounts to 110k GPU hours and is broken down as follows:
@@
  • PFS experiments explore various CLM-to-MLM step ratios (0%–100%, 25%–75%, 50%–50%, 75%–25%, and 100%–0%) across 3 training lengths (12,000, 22,000, and 42,000 steps), focusing on the 610M model size with a 40% masking ratio for MLM. To minimize redundant computation, we reuse checkpoints from the base pretraining runs whenever possible, applying learning rate decay only during the final 2,000 steps. In total, these runs result in 9 new models and account for 112,000 additional training steps and 16k GPU hours.
@@
  • Model evaluation involves fine-tuning a total of 30 final checkpoints: 5 models with 210M parameters, 5 with 1B, and 20 with 610M. In addition, we include 6 intermediate 610M checkpoints from CLM and MLM pretraining. Each model is fine-tuned on 12 datasets, across 6 learning rates and 5 random seeds, resulting in 15,120 fine-tuning runs and consuming 9k GPU hours.
@@ 3 Pretraining with CLM or MLM
As shown in Figure 4, CLM outperforms MLM in downstream performance during the early phase of training (up to around step 5,000). However, as training progresses, MLM models tend to catch up and often surpass CLM, while CLM shows more limited gains in later steps. This suggests that although CLM saturates earlier, it enables more efficient representation learning in fewer steps. Notably, this makes CLM an appealing option for data-scarce scenarios, such as pretraining on low-resource languages, or simply as a warmup stage before MLM-based encoder training.
@@ 5 Continued Pretraining from CLM and MLM Models
Figure 8: Impact of performing MLM CPT on either CLM- or MLM-pretrained models. CPT is conducted for 22,000 steps on 610M models with a 40% masking ratio, following 42,000 steps of initial pretraining.
@@ 6 Related Work
Learning universal text representations has been a central focus in NLP, with MLM emerging as the dominant pretraining objective following the success of models like BERT (Devlin et al., 2019). MLM-based encoders leverage bidirectional attention to capture rich contextual dependencies, enabling strong performance across a wide range of tasks such as classification, question answering, and named entity recognition. Subsequent work has explored architectural extensions (e.g., RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021)) and more efficient pretraining strategies (Clark et al., 2020). In contrast, decoder-only architectures such as GPT-style models were originally optimized for autoregressive generation (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2023; Almazrouei et al., 2023), making them less naturally suited for text representation tasks.
@@
Several recent studies have investigated the key factors influencing the performance of text representation models. BehnamGhader et al. (2024) propose a general framework for adapting decoder-only models to embedding tasks, accompanied by a series of ablations aimed at isolating the primary contributors to their performance. Wettig et al. (2023) explore the impact of the masking ratio on downstream tasks such as sequence classification and question answering, with a focus on learning dynamics, model scale, and fine-tuning behavior. Building on this line of work, we present a comprehensive analysis of how the choice of learning objective and attention mask during training influences downstream representation quality. By employing a unified and controlled pretraining setup, we ensure fair comparisons and minimize confounding factors, allowing for more precise attribution of the effects observed.
@@ 7 Conclusion and Future Work
Our study questions the established belief that Masked Language Modeling (MLM) is universally optimal for encoder pretraining. Based on large-scale experiments, we provide extensive empirical evidence indicating that encoders should not be pretrained exclusively with MLM objectives. Specifically, we show that decoder-only models trained with Causal Language Modeling (CLM) offer distinct advantages in terms of data efficiency, stability, and performance on certain text representation tasks. Importantly, our experiments demonstrate that adapting CLM models through subsequent MLM training consistently outperforms MLM training only, suggesting a nuanced interaction between both training strategies. We hope our findings and resource release will enable future work in this space.
Future work. Future research could explore extending our findings to Vision-Language Models (VLMs), many of which are decoder-based, potentially enhancing their representation learning capabilities. Additionally, future efforts could leverage our insights to pretrain state-of-the-art encoders, building on the combined strengths of CLM and MLM strategies identified in this study.
Limitations. Our study primarily focuses on large-scale ablations conducted exclusively in English. It remains uncertain whether our findings generalize to other languages, as linguistic and structural variations could influence the effectiveness of pretraining strategies. Additionally, our analysis does not account for the common practice of post-training decoder models on specialized datasets, such as those targeting reasoning or mathematical tasks. As such, we have not explored how this additional post-training phase might influence model behaviors or scaling laws relative to models trained from scratch using identical encoder data distributions. Although we observed notable trends up to the 1B parameter scale, future research should investigate whether these dynamics persist, diminish, or evolve at significantly larger model scales and training datasets, particularly considering post-training effects. Finally, in this work, we primarily focus on traditionally reported encoder evaluation settings. However, newer usage paradigms, such as late interaction (Khattab & Zaharia, 2020) or late chunking techniques (Günther et al., 2024; Conti et al., 2025) which heavily rely on bidirectional attention, may reveal even larger performance disparities.
@@ Acknowledgments
We gratefully acknowledge the ADASTRA supercomputer (CINES) for providing technical support and high-performance computing (HPC) resources through projects C1615122, GDA2401, and CAD16490. We also thank the French government for its support through the France 2030 program, as part of the ArGiMi project. This work was further supported by the European Union’s Horizon Europe Research and Innovation Actions (UTTER, grant 101070631), the DECOLLAGE project (ERC-2022-CoG, grant 101088763), and the Portuguese Recovery and Resilience Plan via project C645008882-00000055 (Center for Responsible AI). Additional support was provided by FCT/MECI through national funds and, when applicable, co-funded EU funds under UID/50008: Instituto de Telecomunicações.

### v2

@@ Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM objective or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 38 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM, achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models, reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
@@ 1 Introduction
In this paper, we conduct a controlled study to investigate the impact of training with MLM and CLM objectives when learning textual representations. We compare models with equal architecture and size, trained on equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively on either objective, and then investigate a two-stage pretraining alternative, where training starts with the CLM objective and is followed by MLM. Additionally, given the availability of existing pretrained models, we investigate a continued pretraining scenario, where models are initialized from checkpoints trained exclusively with MLM or CLM, and then trained with MLM for a fraction of the pretraining steps. In total, we train a total of 38 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs, amounting to 110k MI250X GPU hours. Backed by these large-scale experiments, our results show that:
@@ 2 Experimental Setup
Pretraining is performed with a per-device batch size of 12 samples across 192 GPUs, yielding an effective batch size of 2,373,120 tokens. [^5] Inputs consist of variable-length sequences ranging from 12 tokens up to the maximum of 2,048. We employ a Warmup-Stable-Decay (WSD) learning rate schedule: a 2,000-step warmup phase, followed by 38,000 steps with a constant learning rate of 5e-4, [^6] The optimal learning rate was selected from the range 1e-4 to 2e-3 based on experiments using 10% of the training data. ending with a 2,000-step linear decay phase, for a total of 42,000 training steps.
@@
To ensure a fair comparison, all models are fine-tuned using a consistent protocol. Each model is trained for up to 1,000 steps or one full epoch, whichever comes first, using a batch size of 32. To account for differences in model architecture and task requirements, we perform a grid search over 6 learning rates (1e-5, 2e-5, 5e-5, 1e-4, 2e-4, and 5e-4) for each model-dataset pair, with 10% warmup followed by linear decay. The learning rate yielding the best validation performance is selected. Fine-tuning employs bidirectional attention mask with task-specific loss functions: for SC, we use cross-entropy on mean-pooled token embeddings; TC and QA are trained using token-level cross-entropy; and for IR, we rely on the InfoNCE loss (Oord et al., 2018) with in-batch negatives, using mean pooling. To account for the fine-tuning instability commonly observed in BERT-style models for representation learning (Devlin et al., 2019; Lee et al., 2020; Dodge et al., 2020; Zhang et al., 2020), the entire procedure is repeated across 5 random seeds. Fine-tuning is conducted on the in-domain training set, except for NQ and MLDR, for which training is performed on MS MARCO.
@@
SC is assessed with accuracy, TC and QA with F1 score, and IR with NDCG@10. We report results averaged across seeds, along with 95% confidence intervals.
@@
Drawing reliable conclusions about model design choices typically requires scaling up and repeating experiments, as statistically sound results often require a large number of runs and sufficient training to reach minimal convergence (Hoffmann et al., 2022). We therefore carefully design our experiments to ensure robust and well-supported findings. The default pretraining setup uses 100B tokens, which corresponds to 5 times the optimal data budget proposed by Hoffmann et al. (2022) for decoder models in the 1B-parameter range. Pretraining runs are executed on 24 nodes with 8 GPUs each (192 GPUs total), while fine-tuning is carried out on single GPUs. The total compute budget amounts to 110k GPU hours, broken down as follows:
@@
  • PFS experiments explore various CLM-to-MLM step ratios (0%–100%, 25%–75%, 50%–50%, 75%–25%, and 100%–0%) across 3 training lengths (12,000, 22,000, and 42,000 steps), focusing on the 610M model size with a 40% masking ratio for MLM. This setup is also used to evaluate the effect of masking ratio on models pretrained with CLM followed by MLM. To minimize redundant computation, we reuse checkpoints from the base pretraining runs whenever possible, applying learning rate decay only during the final 2,000 steps. In total, these runs result in 17 new models and account for 120,000 additional training steps and 16k GPU hours.
@@
  • Model evaluation involves fine-tuning 50 model checkpoints in total, including 38 final pretrained checkpoints and 12 intermediate ones. Each model is fine-tuned on 12 datasets, across 6 learning rates and 5 random seeds, resulting in 15,120 fine-tuning runs which total about 9k GPU hours. Allocating substantial compute to the evaluation phase is particularly important in representation learning, where fine-tuning can exhibit high variance and instability (Devlin et al., 2019; Lee et al., 2020; Dodge et al., 2020; Zhang et al., 2020).
@@ 3 Pretraining with CLM or MLM
As shown in Figure 4, CLM consistently outperforms MLM in downstream performance during the early stages of training (up to step 10,000 for SC and QA, 20,000 for IR, and even until the end for TC). However, as training progresses, MLM models tend to catch up and often surpass CLM, while CLM shows more limited gains in later steps. This suggests that although CLM saturates earlier, it enables more efficient representation learning in fewer steps. Notably, this makes CLM an appealing option for data-scarce scenarios, such as pretraining on low-resource languages, or simply as a warmup stage before MLM-based encoder training.
@@ 5 Continued Pretraining from CLM and MLM Models
Figure 8: Impact of performing MLM CPT on either CLM- or MLM-pretrained models (denoted as Base). CPT is conducted for 22,000 steps on 610M models with a 40% masking ratio, following 42,000 steps of initial pretraining.
@@ 6 Related Work
Learning universal text representations has been a central focus in NLP, with MLM emerging as the dominant pretraining objective following the success of models like BERT (Devlin et al., 2019). MLM-based encoders leverage bidirectional attention to capture rich contextual dependencies, enabling strong performance across a wide range of tasks such as classification, question answering, and named entity recognition. Subsequent work has explored architectural extensions (e.g., RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), ModernBert (Warner et al., 2024)) and more efficient pretraining strategies (Clark et al., 2020). In contrast, decoder-only architectures such as GPT-style models were originally optimized for autoregressive generation (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2023; Almazrouei et al., 2023), making them less naturally suited for text representation tasks.
@@
Several recent studies have investigated the key factors influencing the performance of text representation models. BehnamGhader et al. (2024) propose a general framework for adapting decoder-only models to embedding tasks, accompanied by a series of ablations aimed at isolating the primary contributors to their performance. Some papers explore how to reintroduce bidirectional attention in causal decoder models, by repeating sequences twice (Springer et al., 2024), training with custom-shaped attention masks (Raffel et al., 2023) or selectively removing the causal attention mask on parts of the sequence in a post-training phase (Kopiczko et al., 2024; Beyer et al., 2024). Wettig et al. (2023) explore the impact of the masking ratio on downstream tasks such as sequence classification and question answering, with a focus on learning dynamics, model scale, and fine-tuning behavior. Building on this line of work, we present a comprehensive analysis of how the choice of learning objective and attention mask during training influences downstream representation quality. By employing a unified and controlled pretraining setup, we ensure fair comparisons and minimize confounding factors, allowing for more precise attribution of the effects observed.
@@ 7 Conclusion and Future Work
Our study challenges the long-standing assumption that Masked Language Modeling (MLM) is the universally optimal approach for encoder pretraining. Through large-scale experiments at the 100B training token scale, we provide robust empirical evidence that encoders can be trained more data-efficiently without relying solely on MLM objectives. In particular, we demonstrate that decoder-only models trained with Causal Language Modeling (CLM) offer clear advantages in data efficiency, training stability, and performance on specific text representation tasks. Moreover, we show that sequential training (first with CLM followed by MLM) is more effective than MLM alone. These results suggest that the most effective path to high-performing encoder models may involve leveraging pretrained decoder models and continuing with MLM-based training, paving the way for future state of the art encoders at a discounted price. We hope our findings and released resources will support further advancements in this area.
Future work. Future research may extend our findings to Vision-Language Models (VLMs), many of which use decoder-based architectures, potentially improving their representation learning capabilities. Further investigations could also apply our insights to pretrain state-of-the-art encoders by capitalizing on the complementary strengths of CLM and MLM, and leveraging the availability of powerful open-source decoders, as suggested in this work.
Limitations. Our experiments are conducted at the 100B training token scale, over 5 times the compute-optimal ratio suggested for decoders in Hoffmann et al. (2022), but our loss curves suggest that further performance improvements are possible with extended training. In extremely high-compute and high-data regimes, the data-efficiency gains from CLM may diminish as they are counterbalanced by increased training capacity. Additionally, our study focuses on large-scale ablations conducted solely in English. It remains uncertain whether our conclusions generalize to other languages, where linguistic and structural differences may affect pretraining efficacy. We also do not examine the impact of common post-training practices applied to decoder models, such as fine-tuning on reasoning or mathematical datasets. Consequently, we have not analyzed how this post-training phase might influence model behavior or scaling trends relative to models pretrained from scratch with encoder-style data. Lastly, our evaluation centers on standard encoder benchmarks. Emerging use cases, such as late interaction (Khattab & Zaharia, 2020) or late chunking (Günther et al., 2024; Conti et al., 2025), which depend heavily on bidirectional attention, may surface even larger performance gaps.
@@ Acknowledgments
We gratefully acknowledge the ADASTRA supercomputer (CINES) for providing technical support and high-performance computing (HPC) resources through projects C1615122, GDA2401, and CAD16490. We also thank the French government for its support through the France 2030 program, as part of the ArGiMi project. This work was further supported by the European Union’s Horizon Europe Research and Innovation Actions (UTTER, grant 101070631), the DECOLLAGE project (ERC-2022-CoG, grant 101088763), and the Portuguese Recovery and Resilience Plan via project C645008882-00000055 (Center for Responsible AI). Additional support was provided by FCT/MECI through national funds and, when applicable, co-funded EU funds under UID/50008: Instituto de Telecomunicações. Finally, we thank Antoine Chaffin for the insightful discussions.